OpenStudy (anonymous):

Sophia invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of her investment after x years. f(x) = 500(1.05)x What was the average rate of change of the value of Sophia's investment from the second year to the fourth year? 14.13 dollars per year 28.25 dollars per year 50.00 dollars per year 56.50 dollars per year

OpenStudy (mathstudent55):

I assume the x is an exponent. First, find f(4). Then find f(2). Then subtract their values and divide by 2.

OpenStudy (anonymous):

Yes

OpenStudy (wolf1728):

That formula should actually be: Total = Principal * (1 + rate) ^ years and I guess we assume the principal is $500 as mathstudent55 said we must find f(4) which is 500*(1.05)^4 = 607.75 Then find f(2) which is 500*(1.05)^2 = 551.25 607.75 -551.25 = 56.50 for 2 years or 28.25 per year which looks like it's the second answer